# Optimizing Attention with Mirror Descent: Generalized Max-Margin Token Selection

### v3

@@ 5 Conclusion
The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the results reported within this paper. This work was supported in part by MathWorks, the MIT-IBM Watson AI Lab, the MIT-Amazon Science Hub, and the MIT-Google Program for Computing Innovation. The authors also thank Roey Magen for valuable feedback on the joint regularization path analysis.
@@ A.2 Lemma for Analyzing The $\ell_{p}$-Norm
$p]{\sum_{i=1}^{d^{2}}(|w_{i}|^{2})^{p/2}}$ ($p]{\left(\sum_{i=1}^{d^{2}}|w_{i}|^{2}\right)^{p/2}}$
=$p]{\|w\|_{2}^{p}}=\|w\|_{2}.$
@@
$p]{\frac{1}{d^{2}}\sum_{i=1}^{d^{2}}|w_{i}|^{p}}\geq\sqrt{\frac{1}{d^{2}}\sum_{i=1}^{d^{2}}|w_{i}|^{2}},$
@@
$\frac{d}{dx}y(x^{p-1}-y^{p-1})=(p-1)x^{p-2}y\leq(p-1)x^{p-1},$
so as we increase $x$, the $\frac{p-1}{p}x^{p}-\frac{p-1}{p}y^{p}$ grows faster than $y(x^{p-1}-y^{p-1})$, so to prove the right inequality we simply need to prove that it holds equality at $x=y$, which is trivially true. Similarly, the left inequality can be proven in a similar way.
@@
$\frac{d}{dx}qx^{q-1}(x-y)=q(q-1)x^{q-2}(x-y)+qx^{q-1},\text{ and}$
$\frac{d}{dx}qy^{q-1}(x-y)=qy^{q-1}.$
When $q\geq 1$,
$\frac{d}{dx}(x^{q}-y^{q})\leq\frac{d}{dx}qx^{q-1}(x-y),$
so because we have
$x^{q}-y^{q}=qx^{q-1}(x-y)=0$
when $x=y$, then
$x^{q}-y^{q}\leq qx^{q-1}(x-y)$
when $x\geq y\geq 0$ if $q\geq 1$. We can use a similar argument for the $10<q<$ 1 case.
@@ A.3 Lemma for Analyzing ERM Objective and Its Gradient
Under Assumption A, $\mathcal{L}(W)$ is bounded from above by $\mathcal{L}_{max}$ and below by $\mathcal{L}_{min}$ for some dataset-dependent constants $\mathcal{L}_{max}$ and $\mathcal{L}_{min}$ that are finite.
@@
Proof  Using the expression in Lemma 17, since $l^{\prime}$ is bounded in a closed interval (Assumption A), the entries in $\sigma(h_{i})$ is always between 00 and $1$, and $\gamma_{i}^{\top}\sigma(h_{i})$ is bounded, then the entries of $\nabla\mathcal{L}(W)$ is bounded by a dataset-dependent bounded, which directly implies this lemma statement.
In the following lemma, we analyze the behaviors of the ($\ell_{p}$-AttSVM) constraint $(X_{it}-X_{i\tau})^{\top}Wz_{i}$ for all $W\in S_{p,\mu_{0}}(W_{\mathrm{mm}}^{\alpha})$ satisfying $\|W\|_{p,p}=\|W_{\mathrm{mm}}^{\alpha}\|_{p,p}$, the result of which is a generalization of (Tarzanagh et al., 2023, Equation 64) for a general $\ell_{p}$ norm.
@@
$h_{i}:=X_{i}Wz_{i},\leavevmode\nobreak\ \leavevmode\nobreak\ \tilde{h}_{i}:=X_{i}Vz_{i},\leavevmode\nobreak\ \leavevmode\nobreak\ l_{i}^{\prime}:=l^{\prime}(\gamma_{i}^{\top}\sigma(h_{i})),\leavevmode\nobreak\ \leavevmode\nobreak\ \textnormal{and}\leavevmode\nobreak\ \leavevmode\nobreak\ s_{i}=\sigma(h_{i}).$
@@
Just as it was in Lemma 18, $l_{i}^{\prime}$ is bounded for any $i\in[n]$ by some bound that is dataset-dependent. Furthermore, because $l$ is decreasing, $-l^{\prime}$ is always non-negative, so an easier approach is to lower-bound the following for each $i\in[n]$,
@@
Note that using Assumption A, $\bar{L}$ is positive and bounded. Hence, using (18) and (A.3), the term $-\langle\mathcal{L}(W),V\rangle$ is positive when
@@
$\begin{split}\sum_{i=1}^{n}(-l_{i}^{\prime})&\left((1+\pi)\left(\sum_{t\in\mathcal{T}_{i}}(\tilde{h}_{i\alpha_{i}}-\tilde{h}_{it})s_{it}(\gamma_{i\alpha_{i}}-\gamma_{it})-2\Gamma A((1-s_{i\alpha_{i}})^{2}+Q_{i}(W))\right)\right.\\ &\left.-\left(\sum_{t\in\mathcal{T}_{i}}(\bar{h}_{i\alpha_{i}}-\bar{h}_{it})s_{it}(\gamma_{i\alpha_{i}}-\gamma_{it})+2\Gamma A((1-s_{i\alpha_{i}})^{2}+Q_{i}(W))\right)\right),\end{split}$ (20)
@@
In both bounds, the main inequality derivation stems from the fact that $\bar{h}_{i\alpha_{i}}>\bar{h}_{i\tau}$ for all $\tau\in[T]$, which we obtain from Lemma 19. Now, we analyze the left double-summation in Equation (21). To analyze the sum, let $\mathcal{I}$ be the subset of $[n]\times[T]$ that contains all $(i,t)$ such that $t\in\mathcal{T}_{i}$. Furthermore, let
$\mathcal{I}_{1}$ :=$:=\left\{(i,t)\in\mathcal{I}\mid\bar{h}_{i\alpha_{i}}-\bar{h}_{it}\leq 1\right\},$
$\mathcal{I}_{2}$ :=$:=\left\{(i,t)\in\mathcal{I}\mid 1<\bar{h}_{i\alpha_{i}}-\bar{h}_{it}\leq 1+\pi\right\},$
$\mathcal{I}_{3}$ :=$:=\left\{(i,t)\in\mathcal{I}\mid\bar{h}_{i\alpha_{i}}-\bar{h}_{it}>1+\pi\right\}.$
@@
$-\frac{\pi}{T}\min_{i\in\mathcal{I}_{1}}\{\gamma_{i}^{gap}\}e^{-R^{\prime}}\max_{i=1}^{n}\{l_{i}^{\prime}\}.$
@@ A.4 Lemma for Analyzing $\ell_{p}$-AttGD
Next, we show several tools for analyzing the algorithm further and for analyzing the Bregman divergence. The following three specifically are from Sun et al. (2022, Lemma 18, 3) and Azizan and Hassibi (2018), and so the proofs are omitted.
@@
This following is a well-known lemma, so the proof is omitted.
Lemma 26 (Bregman Divergences Cosine Law)
For any $w,w^{\prime},w^{\prime\prime}$ that are all vectors or matrices with the same dimensionalities, we have
@@
The following is adapted from Sun et al. (2022, Equation 12) for the case of our attention model. Our proof is quite similar, except that we use our version of the gradient correlation lemma.
@@
So now, let us analyze the term $\langle\nabla\psi(\bar{W}(K)),\bar{W}_{\mathrm{mm}}^{\alpha}\rangle$ using the inductive hypothesis on $k=0,1,...,K-1$. Lemma 27 tells us that
@@
Since this is true for all $k=0,1,...,K-1$, and since $\|W(k)\|_{p,p}$ is increasing in $k$ from Lemma 22 and 20, we can sum all the above inequalities and get the following,
@@ A.5 Lemma for Analyzing Rate of Convergence
$\|W(k+1)\|_{p,p}-\|W(k)\|_{p,p}\leq\frac{p^{2}\eta\leavevmode\nobreak\ \mathcal{O}(\|W(k)\|_{p,p})}{(p-1)\|W(k)\|_{p,p}^{p-1}}\|\nabla\mathcal{L}(W(k))\|_{1,1}.$
@@
We can deduce from Lemma 22 and 20 that $\|W(k)\|_{p,p}$ is increasing in $k$, and with Lemma 4, we can upperbound the increase. Therefore, to approximate that sum above, we can use an integral approximation by ”integrating” over the $\|W(k)\|_{p,p}$ terms.
@@
Suppose Assumptions A holds. Let $\mu^{\prime}$ be that from Lemma 28 if $\mu=\mu_{0}$, and let $R$ the maximum of the $R_{\mu}$ from 28 and $R_{\delta}$ 20. Let $\{W(k)\}_{k\geq 0}$ be the sequence generated by $\ell_{p}$-AttGD. If the initialization $W(0)$ is in $C_{p,\mu^{\prime},R}(W_{\mathrm{mm}}^{\alpha})$, then with a small enough step size $\eta$, we have the following for each $k\geq 0$,
@@ Appendix B Proof of Theorem 6
Proof  The proof is similar to the proof of (Tarzanagh et al., 2024, Theorem 1). Specifically, we need to show that $f(X)=v^{\top}X^{\top}\sigma(XW)$ satisfies the assumptions of (Tarzanagh et al., 2024, Lemma 14), where the nonlinear head is replaced by the linear term $v$. This holds independently of the choice of algorithm or the attention SVM solution. Thus, we omit the details and refer to the proof of (Tarzanagh et al., 2024, Theorem 1).
@@ Appendix C Proof of Theorem 8
Proof
This is a direct implication of Lemma 30
@@ Appendix F On the Convergence of the $\ell_{p}$ Regularization Path for Joint $W$ and $v$
$(v^{(r)},W^{(R)}):=$ $\operatorname*{arg\,min}_{(v,W)}\leavevmode\nobreak\ \mathcal{L}(v,W)$ ($\ell_{p}$-JointRP)
@@
$\bar{W}_{\mathrm{mm}}$ $:=R\Xi W_{\mathrm{mm}},\qquad\bar{v}_{\mathrm{mm}}:=r\Gamma v_{\mathrm{mm}}$ (32)
@@
$\min_{t\neq\alpha_{1}}\leavevmode\nobreak\ (X_{1\alpha_{1}}-X_{1t})^{\top}W^{(R)}z_{1}\leq\Xi R(1-\gamma).$
@@
$\min_{i\in[n]}\leavevmode\nobreak\ \leavevmode\nobreak\ y_{i}(v^{(r)})^{\top}r_{i}$ $\leq\min_{i\in[n]}\leavevmode\nobreak\ \leavevmode\nobreak\ y_{i}(v^{(r)})^{\top}X_{i1}+\sup_{i\in[n]}|(v^{(r)})^{\top}(X_{it}-X_{i1})|$
@@
Appendix G Implementation Details
@@ G.1 Computational Overhead of Algorithm
When compared to the standard gradient descent algorithm, the mirror descent based algorithms, such as the $\ell_{p}$-MD, $\ell_{p}$-AttGD, and $\ell_{p}$-JointGD have additional computational overhead. We claim that the overhead is linear in the size of the parameters for both time complexity and space complexity. For the analysis, we focus on the $\ell_{p}$-AttGD algorithm, but the same analysis can be applied for the other algorithms.
Let $D=d^{2}$, the number of entries in $W(k)$. The gradient descent update rule entails the calculation of the gradient $\nabla\mathcal{L}(W(k))$, and then updating the parameters by subtracting from it $\eta$ times that gradient, giving us $W(k)-\eta\nabla\mathcal{L}(W(k))$.
In $\ell_{p}$-AttGD, computing all entries of $[W(k)]^{+}$ can be done in the following way:
Algorithm 1 Compute $[W(k)]^{+}$
1:Apply $x\mapsto|x|^{p-1}\operatorname{sign}(x)$ on each entry of $W(k)$ to get $W(k)^{\prime}$
2:$[W(k)]^{+}\leftarrow W(k)^{\prime}-\eta\nabla\mathcal{L}(W(k))$
The first step takes $O(D)$, while the second step’s time complexity would be similar as that of the GD algorithm. For computing $W(k+1)$, we just have to apply the mapping $y\mapsto|y|^{1/(p-1)}\operatorname{sign}(y)$ entry-wise to $[W(k)]^{+}$, which would take $O(D)$ time. In total, we have an $O(D)$ time overhead, and for space complexity, we can see that we only need to hold a constant amount of additional space for the computation of each entry, so it requires $O(D)$ additional space.

### v5

@@ 5 Conclusion and Future Work
Our theoretical guarantees are established for single-head, one-layer attention models—a standard abstraction that balances analytical tractability with capturing the essential nonconvex structure of softmax attention. An important direction for future work is extending these convergence results to multi-head and multi-layer architectures, where each head solves separation problems in different subspaces and stacked layers progressively refine feature representations.
@@
The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing computing resources that have contributed to the results reported within this paper. This work was supported in part by MathWorks, Jane Street, the MIT-IBM Watson AI Lab, the MIT-Amazon Science Hub, and the MIT-Google Program for Computing Innovation. The authors also thank Roey Magen for valuable feedback on the joint regularization path analysis.
@@ A.2 Lemma for Analyzing The $\ell_{p}$-Norm
$\sqrt[p]{\sum_{i=1}^{d^{2}}(|w_{i}|^{2})^{p/2}}$ $\leq\sqrt[p]{\left(\sum_{i=1}^{d^{2}}|w_{i}|^{2}\right)^{p/2}}$
$=\sqrt[p]{\|w\|_{2}^{p}}=\|w\|_{2}.$
@@
$\sqrt[p]{\frac{1}{d^{2}}\sum_{i=1}^{d^{2}}|w_{i}|^{p}}\geq\sqrt{\frac{1}{d^{2}}\sum_{i=1}^{d^{2}}|w_{i}|^{2}},$
@@
$\frac{d}{dx}\bigl(y(x^{p-1}-y^{p-1})\bigr)=(p-1)x^{p-2}y\leq(p-1)x^{p-1}.$
Thus, as $x$ increases, $\frac{p-1}{p}x^{p}-\frac{p-1}{p}y^{p}$ grows at least as fast as $y(x^{p-1}-y^{p-1})$. Since both expressions are equal at $x=y$, the right inequality follows. The left inequality can be proven analogously.
@@
$\frac{d}{dx}\bigl(qx^{q-1}(x-y)\bigr)=q(q-1)x^{q-2}(x-y)+qx^{q-1},$
$\frac{d}{dx}\bigl(qy^{q-1}(x-y)\bigr)=qy^{q-1}.$
When $q\geq 1$, we have $q-1\geq 0$ and $x-y\geq 0$, so
$\frac{d}{dx}(x^{q}-y^{q})\leq\frac{d}{dx}\bigl(qx^{q-1}(x-y)\bigr).$
Since both sides equal zero when $x=y$, the first inequality follows. For $0<q<1$, we have q1<0q-1<0, so $x^{q-1}\leq y^{q-1}$ for $x\geq y>0$, which gives $qx^{q-1}\leq qy^{q-1}$. A similar argument then establishes the second inequality.
@@ A.3 Lemma for Analyzing ERM Objective and Its Gradient
Under Assumption A, $\mathcal{L}(W)$ is bounded from above by $\mathcal{L}_{\max}$ and from below by $\mathcal{L}_{\min}$, where $\mathcal{L}(W)$ denotes the objective of (ERM) with fixed $v$, and $\mathcal{L}_{\max}$ and $\mathcal{L}_{\min}$ are finite dataset-dependent constants.
@@
Proof Using the expression in Lemma 17, since $l^{\prime}$ is bounded in a closed interval (Assumption A), the entries in $\sigma(h_{i})$ is always between 0 and $1$, and $\gamma_{i}^{\top}\sigma(h_{i})$ is bounded, then the entries of $\nabla\mathcal{L}(W)$ is bounded by a dataset-dependent bounded, which directly implies this lemma statement.
In the following lemma, we analyze the behaviors of the ($\ell_{p}$-AttSVM) constraint $(X_{it}-X_{i\tau})^{\top}Wz_{i}$ for all $W\in S_{p,\mu_{0}}(W_{\mathrm{mm}}^{\alpha})$ satisfying $\|W\|_{p,p}=\|W_{\mathrm{mm}}^{\alpha}\|_{p,p}$, the result of which is a generalization of (Tarzanagh et al., 2023, Equation (64)) for a general $\ell_{p}$ norm.
@@
$h_{i}:=X_{i}Wz_{i},\penalty 10000\ \penalty 10000\ \tilde{h}_{i}:=X_{i}Vz_{i},\penalty 10000\ \penalty 10000\ l_{i}^{\prime}:=l^{\prime}(\gamma_{i}^{\top}\sigma(h_{i})),\penalty 10000\ \penalty 10000\ \textnormal{and}\penalty 10000\ \penalty 10000\ s_{i}=\sigma(h_{i}).$
@@
Just as it was in Lemma 18, $l_{i}^{\prime}$ is bounded for any $i\in[n]$ by some bound that is dataset-dependent. Furthermore, because $l$ is decreasing, $l^{\prime}$ is always non-negative, so an easier approach is to lower-bound the following for each $i\in[n]$,
@@
Note that using Assumption A, $\bar{L}$ is positive and bounded. Hence, using (18) and (A.3), the term $\langle\mathcal{L}(W),V\rangle$ is positive when
@@
$\begin{split}\sum_{i=1}^{n}(-l_{i}^{\prime})&\left((1+\pi)\left(\sum_{t\in\mathcal{T}_{i}}(\tilde{h}_{i\alpha_{i}}-\tilde{h}_{it})s_{it}(\gamma_{i\alpha_{i}}-\gamma_{it})-2\Gamma A((1-s_{i\alpha_{i}})^{2}+Q_{i}(W))\right)\right.\\ &\left.-\left(\sum_{t\in\mathcal{T}_{i}}(\bar{h}_{i\alpha_{i}}-\bar{h}_{it})s_{it}(\gamma_{i\alpha_{i}}-\gamma_{it})+2\Gamma A((1-s_{i\alpha_{i}})^{2}+Q_{i}(W))\right)\right).\end{split}$ (20)
@@
In both bounds, the main inequality derivation stems from the fact that $\bar{h}_{i\alpha_{i}}>\bar{h}_{i\tau}$ for all $\tau\in[T]$, which we obtain from Lemma 19. Now, we analyze the left double-summation in Equation (21). To analyze the sum, let $\mathcal{I}$ be the subset of $n]\times[T]$ that contains all $(i,t)$ such that $t\in\mathcal{T}_{i}$. Furthermore, let
$\mathcal{I}_{1}$ $:=\left\{(i,t)\in\mathcal{I}\mid\bar{h}_{i\alpha_{i}}-\bar{h}_{it}\leq 1\right\},$
$\mathcal{I}_{2}$ $:=\left\{(i,t)\in\mathcal{I}\mid 1<\bar{h}_{i\alpha_{i}}-\bar{h}_{it}\leq 1+\pi\right\},$
$\mathcal{I}_{3}$ $:=\left\{(i,t)\in\mathcal{I}\mid\bar{h}_{i\alpha_{i}}-\bar{h}_{it}>1+\pi\right\}.$
@@
$\frac{\pi}{T}\min_{i\in\mathcal{I}_{1}}\{\gamma_{i}^{gap}\}e^{-R^{\prime}}\max_{i=1}^{n}\{l_{i}^{\prime}\}.$
@@ A.4 Lemma for Analyzing $\ell_{p}$-AttGD
Next, we show several tools for analyzing the algorithm further and for analyzing the Bregman divergence. The following four specifically are from Sun et al. (2022), Azizan and Hassibi (2018), and Bregman (1967), and so the proofs are omitted.
@@
Lemma 26
(Bregman, 1967, Bregman Divergences Cosine Law) For any $w,w^{\prime},w^{\prime\prime}$ that are all vectors or matrices with the same dimensionalities, we have
@@
The following is adapted from Sun et al. (2022, Equation (12)) for the case of our attention model. Our proof is quite similar, except that we use our version of the gradient correlation lemma.
@@
So now, let us analyze the term $\langle\nabla\psi(\bar{W}(K)),\bar{W}_{\mathrm{mm}}^{\alpha}\rangle$ using the inductive hypothesis on k=0,1,,K1k=0,1,...,K-1. Lemma 27 tells us that
@@
Since this is true for all k=0,1,,K1k=0,1,...,K-1, and since $\|W(k)\|_{p,p}$ is increasing in $k$ from Lemma 22 and 20, we can sum all the above inequalities and get the following,
@@ A.5 Lemma for Analyzing Rate of Convergence
$\|W(k+1)\|_{p,p}-\|W(k)\|_{p,p}\leq\frac{p^{2}\eta\penalty 10000\ \mathcal{O}(\|W(k)\|_{p,p})}{(p-1)\|W(k)\|_{p,p}^{p-1}}\|\nabla\mathcal{L}(W(k))\|_{1,1}.$
@@
We can deduce from Lemmas 22 and 20 that $\|W(k)\|_{p,p}$ is increasing in $k$, and with Lemma 4, we can upperbound the increase. Therefore, to approximate that sum above, we can use an integral approximation by ”integrating” over the $\|W(k)\|_{p,p}$ terms.
@@
Suppose Assumption A holds. Let $\mu^{\prime}$ be that from Lemma 28 if $\mu=\mu_{0}$, and let $R$ the maximum of the $R_{\mu}$ from Lemma 28 and $R_{\delta}$ from Lemma 20. Let $\{W(k)\}_{k\geq 0}$ be the sequence generated by $\ell_{p}$-AttGD. If the initialization $W(0)$ is in $C_{p,\mu^{\prime},R}(W_{\mathrm{mm}}^{\alpha})$, then with a small enough step size $\eta$, we have the following for each $k\geq 0$,
@@ Appendix B Proof of Theorem 6
Proof The proof is similar to the proof of Tarzanagh et al. (2024, Theorem 1). Specifically, we need to show that $f(X)=v^{\top}X^{\top}\sigma(XW)$ satisfies the assumptions of (Tarzanagh et al., 2024, Lemma 14), where the nonlinear head is replaced by the linear term $v$. This holds independently of the choice of algorithm or the attention SVM solution. Thus, we omit the details and refer to the proof of Tarzanagh et al. (2024, Theorem 1).
@@ Appendix C Proof of Theorem 8
Proof This is a direct implication of Lemma 30.
@@ Appendix F On the Convergence of the $\ell_{p}$ Regularization Path for Joint $W$ and $v$
$(v^{(r)},W^{(R)})$= $\operatorname*{arg\,min}_{(v,W)}\penalty 10000\ \mathcal{L}(v,W)$ ($\ell_{p}$-JointRP)
@@
$\bar{W}_{\mathrm{mm}}$ $:=R\Xi W_{\mathrm{mm}},\qquad\bar{v}_{\mathrm{mm}}:=r\Gamma v_{\mathrm{mm}}.$ (32)
@@
$\min_{t\neq\alpha_{1}}\penalty 10000\ (X_{1\alpha_{1}}-X_{1t})^{\top}W^{(R)}z_{1}\leq\Xi R(1-\gamma).$
@@
$\min_{i\in[n]}\penalty 10000\ \penalty 10000\ y_{i}(v^{(r)})^{\top}r_{i}$ $\leq\min_{i\in[n]}\penalty 10000\ \penalty 10000\ y_{i}(v^{(r)})^{\top}X_{i1}+\sup_{i\in[n]}|(v^{(r)})^{\top}(X_{it}-X_{i1})|$
@@
Appendix G Implementation Details and Additional Experiments
@@ G.1 Computational Overhead of Algorithm
When compared to the standard gradient descent algorithm, the mirror descent based algorithms, such as the $\ell_{p}$-MD, $\ell_{p}$-AttGD, and $\ell_{p}$-JointGD have additional computational overhead. To provide a concrete comparison, we analyze the operation counts for updating a parameter matrix $W(k)$ with $D$ entries under three algorithms: GD, Adam, and $\ell_{p}$-MD. We assume the gradient $\nabla\mathcal{L}(W(k))$ is provided as an oracle, since all three algorithms require gradient computation, and focus solely on the update-rule overhead.
For Adam, the algorithm maintains first and second momentum estimates $M(k)$ and $V(k)$, with update rules
$M(k+1)\leftarrow\beta_{1}M(k)+(1-\beta_{1})\nabla\mathcal{L}(W(k)),\quad V(k+1)\leftarrow\beta_{2}V(k)+(1-\beta_{2})\nabla\mathcal{L}(W(k))^{2},$
$W(k+1)\leftarrow W(k)-\eta\cdot\frac{M(k+1)/(1-\beta_{1}^{k})}{\sqrt{V(k+1)/(1-\beta_{2}^{k})}+\varepsilon}.$
Updating $M(k+1)$ requires two entry-wise multiplications and one addition; updating $V(k+1)$ requires two multiplications, one squaring, and one addition; updating $W(k+1)$ requires three divisions, one subtraction, one addition, one multiplication, and one square root. In total, Adam requires fourteen entry-wise operations per parameter.
For GD, the update is simply
$W(k+1)\leftarrow W(k)-\eta\nabla\mathcal{L}(W(k)),$
requiring only one multiplication and one subtraction, i.e., two entry-wise operations per parameter.
For $\ell_{p}$-MD, the update follows Algorithm $\ell_{p}$-AttGD. Computing $W(k)]^{+}_{ij}$ requires one entry-wise sign operation, one power, one absolute value, two multiplications, and one subtraction. The final computation requires one power, one absolute value, one sign, and one multiplication. In total, $\ell_{p}$-MD requires ten entry-wise operations per parameter.
All of these operations scale linearly with the number of parameters $D$. Thus, despite introducing richer $\ell_{p}$ geometry, $\ell_{p}$-MD remains strictly more efficient than Adam by four operations per parameter, while the overhead relative to GD is modest, linear in the number of parameters, and easily offset by the induced sparsity benefits. This analysis demonstrates that our algorithm achieves a favorable complexity-accuracy tradeoff: its computational overhead is provably negligible and in fact lighter than Adam, while yielding sparsity and interpretability benefits that neither GD nor Adam can provide.